OpenStudy (anonymous):

what's arccsc(-2) equal to

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

Let x = arccsc(-2) x = arccsc(-2) csc(x) = csc(arccsc(-2)) csc(x) = -2 1/sin(x) = -2 1 = -2*sin(x) 1/(-2) = sin(x) -1/2 = sin(x) sin(x) = -1/2 Now use the unit circle to find that sin(-pi/6) = -1/2 So x = -pi/6 This means that arccsc(-2) = -pi/6

OpenStudy (anonymous):

I keep coming up with 11pi/6 but am told it is -pi/6. Why would i choose one over the other?

OpenStudy (anonymous):

read what i said and try :)

OpenStudy (anonymous):

I would think that sin(x)=-1/2; then x=11pi/6. I understand that pi/6 is the co terminal, but why would I choose it rather than 11pi/6? sorry, but i am confused

OpenStudy (anonymous):

it's -30 degrees

OpenStudy (anonymous):

yes, since -pi/6 = -30 degrees

OpenStudy (anonymous):

Well 330 is a coterminal angle to -30 degrees

OpenStudy (anonymous):

I see...thank you

OpenStudy (anonymous):

medal ^^